# Binary-Report Peer Prediction for Real-Valued Signal Spaces

### v1

@@ Abstract
Theoretical guarantees about peer prediction mechanisms typically rely on the discreteness of the signal and report space. However, we posit that a discrete signal model is not realistic: in practice, agents observe richer information and map their signals to a discrete report. In this paper, we formalize a model with real-valued signals and binary reports. We study a natural class of symmetric strategies where agents map their information to a binary value according to a single real-valued threshold. We characterize equilibria for several well-known peer prediction mechanisms which are known to be truthful under the binary report model. In general, even when every threshold would correspond to a truthful equilibrium in the binary signal model, only certain thresholds remain equilibria in our model. Furthermore, by studying the dynamics of this threshold, we find that some of these equilibria are unstable. These results suggest important limitations for the deployment of existing peer prediction mechanisms in practice.
@@ 1 Introduction
In practice, however, the signal space of information that agents receive is much richer than the small number of categories typically studied in the literature. An essay contains far more information than a simple $H$ or $L$ impression. As a result, one would expect the same agent to form a range of posterior beliefs based on these more nuanced impressions. What are the implications of this more nuanced signal space? [^1] One implication which we do not focus on is the well-known issue of spurious correlation, such as adopting a strategy of conditioning one’s report on the first word of the essay; see § 7. Existing results often rely on a clear separation in the joint distribution $P$ to control the incentives, making it unclear how one would expect these mechanisms to perform in practice.
To address this question, we introduce and study a model of binary-report peer prediction for real-valued signal spaces. In this model, agents receive a real number as their signal and then based on it select a report of $H$ or $L$. We focus on the natural class of threshold strategies, where an agent reports $H$ if and only if their signal exceeds a particular threshold. For example, an agent may report $H$ if they deem the quality of an essay to be above 0.7, and $L$ otherwise.
We generally find that mechanisms which are truthful under the original binary signal model fail to yield correct incentives under the more nuanced model with real-valued signals. Intuitively, agents with signals near the threshold may have an incentive to misreport, as the separation that enforces truthfulness breaks down. We give necessary and sufficient conditions for a threshold to be an equilibrium, i.e., not suffer from this issue, in a variety of peer prediction mechanisms. Furthermore we study dynamics arising if a small fraction of agents with the greatest incentive to do so change their report, causing the threshold to slowly shift. We then distinguish stable equilibria under our dynamics as reasonable to arise naturally, from unstable equilibria.
@@ 1.1 Motivating example: Output Agreement (OA)
Consider Output Agreement (OA), which provides a reward of 1 if an agent’s report agrees with a peer and 0 otherwise [Von Ahn and Dabbish, 2004, 2008]. In our more nuanced, real-valued signal model, each agent $i$ recieves a signal we model as a random variable $X_{i}=Z+Z_{i}$ where $Z$ and $Z_{i}$ are both unit normals with $Z$ shared by all agents and $Z_{i}$ unique to agent $i$. In a peer grading setting the signals $X_{i}$ might capture a fine-grained assessment of the quality of an assignment.
@@
In contrast, the traditional signal structure for OA is usually taken to be finite, so take the simplest binary case where the set of signals is $\{H,L\}$. The mechanism designer has some intended mapping between the real-valued and binary signal spaces. Given the interpretation of $H$ and $L$ as “high” and “low”, it is natural for this mapping to be determined by a real-valued threshold $\tau$, with all values above the threshold corresponding to $H$ and all below it corresponding to $L$. In peer grading this naturally corresponds to an intution that a “good-enough” assignment might be graded as satisfactory while a poor one would be deemed unsatisfactory.
@@
Let us suppose first that the mechanism designer announces a desired threshold of $\tau=0$. In the binary signal model with this threshold and distribution (i.e. agents only receive a signal of $H$ or $L$ as determined by $\tau$), OA is truthful because agent $i$ receiving $H$ makes it more likely that all other agents received $H$ than $L$ and vice versa ($\Pr[H|H],\Pr[L|L]>0.5$).
@@
In the real-valued case, we can no longer speak of OA as truthful because it is no longer a direct revelation mechanism, but we can ask whether reporting as intended by $\tau$ is a (Bayes-Nash) equilibrium. It turns out $\tau=0$ is such an equilibrium. For example, if $x_{i}=0.11$, we have $\Pr[X_{j}>0|x_{i}]\approx 0.52>0.5$, so reporting $H$ as intended is indeed the best response.
@@
Now suppose the desired threshold were $\tau=0.1$. In the binary signal model, $\Pr[H|H]=\Pr[X_{j}>0.1|X_{i}>0.1]\approx 0.64$ and $\Pr[L|L]\approx 0.69$ are both greater than $0.5$, so OA remains truthful. Now in the real-valued model consider an agent receiving a signal of $x_{i}=0.11$. As $x_{i}$ is just above the threshold, the intended report is $H$. But $\Pr[X_{j}>0.1|x_{i}]\approx 0.48<0.5$. Thus the best response is $L$, so reporting according to $\tau$ is not an equilibrium. The above analysis will give the same qualitative result for any finite nonzero $\tau$, intuitively because with a positive threshold any agent receiving a signal slightly above the threshold will think that the average agent will have a signal below it and vice versa for negative thresholds. Thus, the only finite initial threshold $\tau$ for which reporting according to $\tau$ is an equilibrium is $\tau=0$.
@@
Given that $\tau=0.1$ is not an equilibrium, what should we expect to happen if the mechanism designer announces it? We saw that agents with signals $x_{i}$ slightly above $0.1$ have an incentive to report $L$ instead of $H$. This incentive weakens as $x_{i}$ grows, so it is natural to expect only a small fraction of agents to deviate from the intended strategy, specifically those with signals closest to but above $0.1$. But effectively, this raises our threshold from $\tau_{0}=0.1$ to some higher value, say $\tau_{1}=0.13$. Since the same argument applies to our new threshold, in future tasks further deviations could lead to $\tau_{2}=0.16$. Thus, we should expect that, over time, the system will converge toward the “trival” equilibrium at $\tau=\infty$ where agents report $L$ regardless of their signal. The same logic shows that announcing $\tau<0$ will lead toward the trivial “always $H$” equilibrium at $\tau=-\infty$.
@@
From a dynamical systems perspective, the above shows that the $\tau=0$ equilibrium is unstable and can only be realized by starting exactly at it. Since there are no other finite equilibria, any variation or small error in the model will inexorably lead to a threshold that tends toward infinity. Given the inherent noisiness of many peer prediction applications, we argue similar to Shnayder et al. [2016] that the only reasonable equilibra in practice are stable ones. Stability is thus a crucial equilibrium refinement for peer prediction.
@@ 1.2 Results
[…]
In the binary signal setting, a multi-task mechanism proposed by Dasgupta and Ghosh [2013] (DG) fixes the uninformative equilibria problem in OA. The DG mechanism adds a penalty corresponding to the empirical frequency of agents’ reports across other tasks, effectively discouraging blind agreement and leading to strictly worse payments in uninformative equilibria. Under our richer signal model, we provide necessary and sufficient conditions for a threshold to be an equilibrium. Under some conditions like monotonicity, these equilibria are stable, while the uniformative equilibria are unstable. We conclude that DG is potentially useful but may only allow limited control over the equilibrium threshold, i.e., what $H$ and $L$ end up meaning.
We perform similar analyses for two other more complex mechanisms, Determininant Mutual Information (DMI) [Kong, 2020, 2024] and the Robust Bayesian Truth Serum (RBTS) [Witkowski and Parkes, 2012]. Under DMI, an interesting consequence of our model is the benefit of coordinating reports across tasks. For all four mechanisms we study we augment our general analysis with a worked example of the Gaussian case and numerical illustrations of more complex distributions that exhibit skewness and multimodality.
In general, our results show that realistic behavior under common peer prediction mechanisms departs significantly from the traditional discrete-signal theory. With a richer signal space, agents will essentially redefine the meaning of the binary $H$ and $L$ signals based on community norms. This paper therefore serves as a launching point for uncovering how behavior changes under peer prediction in the real world. In particular, our results will help practitioners better understand the implications of their choice of mechanism.
@@ 1.3 Related Work
For a general overview of peer prediction, we direct the reader to Faltings [2023], Lehmann [2024], Frongillo and Waggoner [2024]. Most relevant to our work is the distinction between peer prediction mechanisms which are minimal, in that agents only report their signal (as in OA), and those where additional information about agent beliefs is reported (e.g. the agent’s posterior belief about the report of another agent as in the Bayesian Truth Serum [Prelec, 2004] and its variants). Another important dimension is whether the mechanism is for a single task or a larger mult-task collection where reports from unrelated tasks can be compared to improve incentives. We examine two multi-task mechanisms (DG and DMI) and a non-minimal one (RBTS). While more robust than OA, our results show that neither of these structures avoids the issues our work raises.
[…]
Looking closer, two key differences between our work and Shnayder et al. [2016] are the signal model and strategic model. For the signal model, they consider a discrete signal space matching the report space, as usual. For the strategic model, they consider agents learning whether to switch to another strategy in a discrete set, whereas we allow agents to choose any threshold strategy. As a result, our results differ substantially in some cases: for OA, the truthful equilibrium is unstable even in a technical sense (measure zero basin of attraction), and for DG, “truthfulness” as given by the mechanism designer’s desired threshold need not be an equilibrium.
We note that the model and tools in our paper also bear some resemblance to past work in the epistemic democracy literature. Specifically, Duggan and Martinelli [2001] and Meirowitz [2002] study a common-value voting setting, where there is a ground truth “correct” alternative that all agents prefer, and they each receive signals about which alternative that is. As in our paper, the alternative space is binary while the signal space is continuous. The authors similarly identify threshold equilibria, where agents map their signals to a vote for either alternative according to a cutoff point, and prove that under some conditions on the continuity and monotonicity of the prior densities over signals in each world state, there is a unique symmetric Bayes-Nash equilibrium characterized by a threshold. We note that though the tools are similar (equilibria are characterized by evaluating conditional probabilities at the threshold), the setting and goals of the papers are quite different. (The authors aim to show that the “right” alternative is chosen with high probability over the limit of the number of agents.)
@@ 2 Output Agreement
We begin by analyzing the popular peer prediction mechanism Output Agreement (OA). Output Agreement is a minimal, single task mechanism: each agent $i$ submits a report for a task and receives a positive payment if their report matches that of another agent $j$ on the same task. In the binary signal model, truthful reporting forms a correlated equilibrium under some assumptions about the information structure. Specifically, truthfulness holds under strong diagonalization, where the conditional probability of another signal is maximized by the signal being conditioned on ($\Pr[H\mid H]>\Pr[L\mid H]$ and $\Pr[L\mid L]>\Pr[H\mid L]$). However, uninformative equilibria where agents misreport their information also exist. Specifically, agents can coordinate to all submit the same report and each receive a maximum payment. Nonetheless, Output Agreement remains popular because of its simplicity. We thus aim to characterize OA under our richer signal space model to help practitioners understand how behavior occurs in the real world.
@@ 2.1 Model
We begin by introducing the model that, with minor variations to accommodate the form of different peer prediction mechanisms, will be used througout the paper. There are $n$ agents; each agent $i$ receives a signal $X_{i}\in\mathbb{R}$ representing information gained about a shared task. The signals are drawn from a joint distribution $\mathcal{D}$ that is symmetric in the sense that each agent $i$ has (1) the same marginal $F(x)=\Pr[X_{i}=x]$, and (2) the same posterior distribution $\beta(x)=\Pr[X^{\prime}=\cdot\mid X_{i}=x]$. Identical marginals are realistic when the signals are exchangeable, e.g. when they have the same conditional distribution over some latent variable $\theta$.
Let the report space be =$\mathcal{R}=\{L,H\}$. Each agent $i$ calculates a report $r_{i}\in\mathcal{R}$ for the task according to a deterministic strategy $\sigma_{i}:\mathbb{R}\to\mathcal{R}$ mapping from signal to report space. The Output Agreement mechanism pays each agent $i$ an amount $M_{\textrm{OA}}(r_{i},r_{j})$ as a function of reports $r_{i},r_{j}$, where
@@
Thus the (ex-interim) expected utility for playing strategy $\sigma_{i}$ when $j$ plays according to $\sigma_{j}$ is

### v2

# Peer Prediction with More Signals than Reports

@@ Abstract
Peer prediction mechanisms are typically proposed and analyzed under the assumption that the report and signal spaces are identical. In practice, however, agents often observe richer information which they then map to a coarser report space. Motivated by this discrepancy between theory and practice, we initiate the study of peer prediction mechanisms with signal spaces that are richer than the report space. We begin by formalizing a model with real-valued signals and binary reports. In this setting, it is natural to study symmetric threshold strategies, where agents map their signals to binary reports according to a single real-valued threshold. For several well-known binary-report peer prediction mechanisms, we show that most equilibria under the original assumption of binary signals are no longer equilibria in our model. Furthermore, dynamic analysis proves that some of the remaining thresholds are unstable. These results extend beyond real-valued signals and binary reports to settings where the signal space is finer-grained than the report space.
While the results above suggest important limitations for the deployment of existing peer prediction mechanisms in practice, we also use them to develop a new, more robust mechanism. This mechanism generates a larger number of stable threshold equilibria under our model, thus allowing the designer more flexibility in choosing how agents map their signals to reports.
@@ 1 Introduction
In practice, however, the signal space of information that agents receive is much richer than the small number of categories typically studied in the literature. An essay contains far more information than a simple $H$ or $L$ impression. [^1] One implication of a more nuanced signal space which we do not focus on is the well-known issue of spurious correlation, such as adopting a strategy of conditioning one’s report on the first word of the essay; see § 6. Existing results often rely on the assumption of a one-to-one mapping from signals to reports, making it unclear how one would expect these mechanisms to perform in practice.
To address this question, we introduce and study a model of peer prediction with richer signal spaces. We primarily focus on the binary-report setting, where a designer wishes to elicit binary information (e.g. is an essay’s equality $H$ or $L$?) from an agent about a task. Here we assume that agents receive a real number as their signal and then based on it select a report of $H$ or $L$. This real-valued model captures settings where agents form a range of posterior beliefs about a task. We focus on a natural class of threshold strategies, where an agent reports $H$ if and only if their signal exceeds a particular threshold. For example, an agent may report $H$ if they deem the quality of an essay to be above 0.7, and $L$ otherwise.
We generally find that mechanisms which are truthful under the original binary signal model fail to yield correct incentives under the more nuanced model with real-valued signals. Intuitively, agents with signals near the threshold may have an incentive to misreport, as the separation that enforces truthfulness breaks down. We give necessary and sufficient conditions for a threshold to be an equilibrium, i.e., not suffer from this issue, in a variety of peer prediction mechanisms. Furthermore we study dynamics arising if a small fraction of agents with the greatest incentive to do so change their report, causing the threshold to slowly shift. We then distinguish stable equilibria under our dynamics as reasonable to arise naturally, in contrast to unstable equilibria.
@@ 1.1 Motivating example: Output Agreement (OA)
Consider Output Agreement (OA), which provides a reward of 1 if an agent’s binary report of $L$ or $H$ agrees with a peer and 0 otherwise [Von Ahn and Dabbish, 2004, 2008]. Traditional analysis assumes the signal space is the same as the report space, i.e. $\{L,H\}$. In contrast, our real-valued signal model assumes each agent $i$ recieves a real-valued signal $X_{i}$ from some joint distribution. As a running example, let $X_{i}=Z+Z_{i}$ with $Z$ and $Z_{i}$ both unit normals, where $Z$ is shared by all agents and $Z_{i}$ is signal noise unique to agent $i$. In a peer grading setting the signals $X_{i}$ might capture a fine-grained assessment of the quality of an assignment.
@@
The mechanism designer has some intended mapping between the real-valued signals and binary reports. Given the interpretation of $H$ and $L$ as “high” and “low”, it is natural for this mapping to have a real-valued threshold $\tau$, with values above the threshold mapping to $H$, and below mapping to $L$. In peer grading this assumption naturally corresponds to an intution that a “good-enough” assignment might be graded as satisfactory while a poor one would be deemed unsatisfactory.
@@
Let us suppose first that the mechanism designer announces a desired threshold of $\tau=0$. In the binary signal model with this threshold and distribution (i.e. agents only receive a signal of $H$ or $L$ as determined by $\tau$), OA is truthful: after receiving $H$, an agent believes it is more likely that their peer received $H$ than $L$ and vice versa ($\Pr[H\mid H],\Pr[L\mid L]>0.5$).
@@
In the real-valued case, we can no longer speak of OA as truthful because it is no longer a direct revelation mechanism, but we can ask whether reporting as intended by $\tau$ is a Bayes-Nash equilibrium. It turns out $\tau=0$ is such an equilibrium. For example, if $x_{i}=0.11$, we have $\Pr[X_{j}>0\mid x_{i}]\approx 0.52>0.5$, so reporting $H$ as intended is indeed the best response.
@@
Now suppose the desired threshold were $\tau=0.1$. In the binary signal model, $\Pr[H\mid H]=\Pr[X_{j}>0.1\mid X_{i}>0.1]\approx 0.64$ and $\Pr[L\mid L]\approx 0.69$ are both greater than $0.5$, so OA remains truthful. Now in the real-valued model consider an agent receiving a signal of $x_{i}=0.11$. As $x_{i}$ is just above the threshold, the intended report is $H$. But $\Pr[X_{j}>0.1\mid x_{i}]\approx 0.48<0.5$. Thus the best response is $L$, so reporting according to $\tau$ is not an equilibrium. The above analysis will give the same qualitative result for any finite nonzero $\tau$, intuitively because for $\tau>0$ an agent whose signal lies just above the threshold still assigns higher probability to a peer’s signal falling below $\tau$ than above it; analagous reasoning applies for negative thresholds. Thus, the only finite initial threshold $\tau$ for which reporting according to $\tau$ is an equilibrium is $\tau=0$.
@@
Given that $\tau=0.1$ is not an equilibrium, what should we expect to happen if the mechanism designer announces it? We saw that agents with signals $x_{i}$ slightly above $0.1$ have an incentive to report $L$ instead of $H$. This incentive weakens as $x_{i}$ grows, so it is natural to expect only a small fraction of agents to deviate from the intended strategy, specifically those with signals closest to but above $0.1$. But effectively, this raises our threshold from $\tau_{0}=0.1$ to some higher value, say $\tau_{1}=0.13$. Since the same argument applies to our new threshold, in future tasks further deviations could lead to $\tau_{2}=0.16$. Thus, we should expect that, over time, the system will converge toward the “trival” equilibrium at $\tau=\infty$ where agents report $L$ regardless of their signal. The same logic shows that announcing $\tau<0$ will lead toward the trivial “always HH” equilibrium at $\tau=$-$\infty$.
@@
From a dynamical systems perspective, the above shows that the $\tau=0$ equilibrium is unstable and can only be realized by starting exactly at it. As there are no other finite equilibria, any variation or small error in the model will send the threshold toward infinity. Given the inherent noisiness of many peer prediction applications, we argue similar to Shnayder et al. [2016b] that the only reasonable equilibra in practice are stable ones. Stability is thus a crucial equilibrium refinement for peer prediction.
@@ 1.2 Results
We explore the generality of our observation that moving to a real-valued signal model substantially changes the incentives and behavior of peer prediction mechanisms. We begin in § 2.1 by formally analyzing the Output Agreement (OA) mechanism mentioned above. As discussed, not only do uninformative equilibria exist where all agents submit the same report, but they also remain stable under dynamics. That is, at least for some natural distributions, we would expect any initial threshold to drift toward $\pm\infty$ over time, eventually leading to blind agreement.
In the binary signal setting, a multi-task mechanism proposed by Dasgupta and Ghosh [2013] (DG) fixes the uninformative equilibria problem in OA. The DG mechanism adds a penalty corresponding to the empirical frequency of agents’ reports across other tasks, effectively discouraging blind agreement and leading to strictly worse payments in uninformative equilibria. Under our richer signal model, we provide necessary and sufficient conditions for a finite threshold to be an equilibrium. Under some conditions like monotonicity, these equilibria are stable, while the uniformative equilibria are unstable. However, we find that in many cases there exists only one such equilibrium determined by the underlying information structure, meaning the designer has no control over how agents decide to map their signals to $H$ versus $L$. We conclude that DG is potentially more useful than OA, but only allows limited control over the equilibrium threshold.
We perform similar analyses for two other more complex mechanisms, the Robust Bayesian Truth Serum (RBTS) [Witkowski and Parkes, 2012] and Determininant Mutual Information (DMI) [Kong, 2020, 2024]. In both cases, similar characterizations of equilibria and stability emerge as DG. For all four mechanisms we augment our general analysis with a worked example of the Gaussian case and numerical illustrations that exhibit similar results under skewness and multimodality. Finally, we demonstrate that these limitations persist in settings of finer-grained signals beyond the real-valued case, as well as in non-binary report settings using the Correlated Agreement (CA) mechanism [Shnayder et al., 2016a]. Collectively, then, we find that current peer prediction mechanisms are inflexible under a richer signal space in many informational settings.
A natural next question, then, is whether we can come up with a new framework that grants the designer more choice in how agents map information to reports. In Section 5, we propose the following report-mapping framework: ask agents for a raw, non-binary report using the Correlated Agreement mechanism, e.g., agents submit $\hat{L}$, $\hat{M}$ (medium), and $\hat{H}$. Then the designer can impose a mapping back to a final binary report of $L$ or $H$, with potentially more flexibility. Using our results for non-binary reports, we show that CA parameters can be tuned within this mapping framework to robustly generate a larger number of stable thresholds than binary-report mechanisms can achieve under the Gaussian model.
Our paper serves as a launching point for uncovering how behavior changes under peer prediction in the real world. We show that under a more realistic signal model, the designer has much less flexibility to establish a mapping between information and reports under traditional binary-report mechanisms. We also show that some of this flexibility can be recovered by leveraging the Correlated Agreement mechanism and mapping larger report spaces back to a binary decision. Our results clarify the implications of choosing various mechanisms in practice, and provide tools to better control the semantics of their reports.
@@ 1.3 Related Work
For a general overview of peer prediction, we direct the reader to Faltings [2023], Lehmann [2024], Frongillo and Waggoner [2024]. Most relevant to our work is the distinction between peer prediction mechanisms which are minimal, in that agents only report their signal (as in OA), and those where additional information about agent beliefs is reported (e.g. the agent’s posterior belief about the report of another agent as in the Bayesian Truth Serum [Prelec, 2004] and its variants). Another important dimension is whether the mechanism is for a single task or a larger multi-task collection where reports from unrelated tasks can be compared to improve incentives. We examine two multi-task mechanisms (DG and DMI) and a non-minimal one (RBTS). While more robust than OA, our results show that neither of these structures avoids the issues our work raises.
To our knowledge, our approach of assuming agents have a fundamentally richer signal than they are asked to reveal is novel. The idea of using dynamical stability for equilibrium refinement in peer prediction was introduced by Shnayder et al. [2016b]. That work looks at a population of agents playing strategies from a discrete set, such as $\{$ truthful, always $H$, always $L$ }\}. Using replicator dynamics, they measure the stability of equilibria by their basin of attraction, the volume of initial conditions leading to that equilbrium. At a very high level, their conclusions bear similarity to ours: useful equilibria are less robust for OA, and multi-task mechanisms like DG are more robust.
Looking closer, two key differences between our work and Shnayder et al. [2016b] are the signal model and strategic model. For the signal model, they consider a discrete signal space matching the report space, as usual. For the strategic model, they consider agents learning whether to switch to another strategy in a discrete set, whereas we allow agents to choose any threshold strategy. As a result, our results differ substantially in some cases: for OA, the truthful equilibrium is unstable even in a technical sense (measure zero basin of attraction), and for DG, “truthfulness” as given by the mechanism designer’s desired threshold need not be an equilibrium.
The model and tools in our paper also bear some resemblance to work in the epistemic democracy literature. Specifically, Duggan and Martinelli [2001] and Meirowitz [2002] study a common-value voting setting, where there is a ground truth “correct” alternative that all agents prefer, and they each receive signals about which alternative that is. As in our paper, the alternative space is binary while the signal space is continuous. The authors similarly identify threshold equilibria, where agents map their signals to a vote for either alternative according to a cutoff point, and prove that under continuity and monotonicity conditions over the prior densities, there is a unique symmetric Bayes-Nash equilibrium characterized by a threshold. We note that though the tools are similar, the setting and goals of the papers are quite different: the authors aim to show that the “right” alternative is chosen with high probability as the number of agents grows large.
@@ 2 Output Agreement
We begin by analyzing the popular peer prediction mechanism Output Agreement (OA) when signals are real-valued and reports are binary. Output Agreement is a minimal, single task mechanism: each agent $i$ submits a report for a task and receives a positive payment if their report matches that of another agent $j$ on the same task. In the binary signal model, truthful reporting forms a Bayes-Nash equilibrium under some assumptions about the information structure. Specifically, truthfulness holds under strong diagonalization, where the conditional probability of another signal is maximized by the signal being conditioned on ($\Pr[H\mid H]>\Pr[L\mid H]$ and $\Pr[L\mid L]>\Pr[H\mid L]$). However, uninformative equilibria where agents misreport their information also exist. Specifically, agents can coordinate to all submit the same report and each receive a maximum payment. Nonetheless, Output Agreement remains popular because of its simplicity. We characterize OA under our richer signal space model to help practitioners understand how behavior occurs in the real world.
@@ 2.1 Model
We begin by introducing the model that, with minor variations to accommodate the form of different peer prediction mechanisms, will be used througout the paper. There are $n$ agents; each agent $i$ receives a signal $X_{i}\in\mathbb{R}$ representing information gained about a shared task. The signals are drawn from a joint distribution $\mathcal{D}$ that is symmetric in the sense that each agent $i$ has (1) the same marginal with CDF $F(x)=\Pr[X_{i}\leq x]$, and (2) the same posterior distribution $\beta(x)=\Pr[X^{\prime}=\cdot\mid X_{i}=x]$. Identical marginals are realistic when the signals are exchangeable, e.g. when they have the same conditional distribution over some latent variable $\theta$.
Let the report space be $\mathcal{R}=\{L,H\}$. Each agent $i$ calculates a report $r_{i}\in\mathcal{R}$ for the task according to a deterministic strategy $\sigma_{i}:\mathbb{R}\to\mathcal{R}$ mapping from signal to report space. The Output Agreement mechanism pays each agent $i$ an amount $M_{\textrm{OA}}(r_{i},r_{j})$ as a function of reports $r_{i},r_{j}$, where
@@
Thus the (interim) expected utility for playing strategy $\sigma_{i}$ when $j$ plays according to $\sigma_{j}$ is